OpenStudy (anonymous):

what is 5cis(π) in rectangular form?

OpenStudy (ikram002p):

cis ?

OpenStudy (anonymous):

Its a complex numbers question. cis(a) = cos(a) +isin(a) so 5cis(pi) = 5(cos(pi)+isin(pi)) =-5 However this solution looked at bit dodgy so I want someone to confirm that its correct.

zepdrix (zepdrix):

\[\Large\rm 5cis(\pi)=5\left[\cos(\pi)+\mathcal i \sin(\pi)\right]\]\[\Large\rm 5cis(\pi)=5\left[-1+\mathcal i(0)\right]\]\[\Large\rm 5cis(\pi)=-5+0\mathcal i\]Seems right, yah? :)

OpenStudy (anonymous):

He is right.

OpenStudy (anonymous):

Cool thanks for the replies!